the area of asquare park is (x2 + 8x +16) square units. find its sides by finding its factors

Found 2 solutions by Fombitz, ewatrrr:
Answer by Fombitz(32388)
Answer by ewatrrr(24785)

A = (x2 + 8x +16) = (x+4)(x+4) = s•s, s = (x+4)